# Robotic Sorting Systems: Robot Management and Layout Design Optimization

### v3

@@ Abstract
In the contemporary logistics industry, automation plays a pivotal role in enhancing production efficiency and expanding industrial scale. In particular, autonomous mobile robots have become integral to modernization efforts in warehouses. One noteworthy application in robotic warehousing is the robotic sorting system (RSS), which is distinguished by its cost-effectiveness, simplicity, scalability, and adaptable throughput control. Previous research on RSS efficiency often assumed an ideal robot management system, ignoring potential traffic delays and assuming constant travel times. We introduce a novel robot traffic management method, named Rhythmic Control for Sorting Scenario (RC-S), for RSS operations, along with an estimation formula that establishes the quantitative relationship between system performance and configurations. Simulations validate that RC-S reduces average service time by 10.3% compared to the classical cooperative A* algorithm, while also improving throughput and runtime. Based on the performance analysis of RC-S, we develop a layout optimization model that considers system configurations, desired throughput, and costs to minimize expenses and determine the optimal layout. Numerical studies show that facility costs dominate at lower throughput levels, while labor costs prevail at higher throughput levels. Additionally, due to traffic efficiency limitations, RSS is well-suited for small-scale operations like end-of-supply-chain distribution centers.
@@ 1 Introduction
Improved logistics and delivery services have fueled the rapid growth of e-commerce in the 21st century. The development of stable supply chains, same-day or next-day delivery options, and hassle-free return policies increase consumer confidence in online shopping (BusinessWire,, 2023). In order to handle the increasing volume of online orders, a new generation of warehouses specifically catering to individual customers has become a hot topic for logistics companies. This kind of warehouse efficiently meets the demand for small orders with tight delivery schedules through the implementation of automated equipments (Boysen et al., 2019b, ). For instance, robotic mobile fulfill system (RMFS), which employs a rack-moving mechanism, is widely used in intra-warehouse logistics, such as Amazon’s KIVA systems (Wurman et al.,, 2008).
The main activities performed in warehouse include: (1) receiving, (2) transfer and put away, (3) order picking/selection, (4) accumulation/sorting, (5) cross-docking, and (6) shipping (De Koster et al.,, 2007). In this paper, we focus on the sorting process. Sorting involves categorizing and consolidating parcels according to their order information and shipping destinations. Conventional sorting systems commonly employ conveyor-based sorters , where the actuators move along with the conveyor belt, sequentially passing through each outbound station and releasing the loaded parcels at the appropriate locations (Boysen et al., 2019a, ). These systems are highly appreciated for its efficiency and stability, while suffering from the inflexibility and significant space occupation (Boysen et al.,, 2023).
@@
A novel sorting system utilizing mobile robots attracts increasing attention in recent years, which is called the robotic sorting system (RSS) (Zou et al.,, 2021). Figure 1(a) presents a dual-layer solution of RSS. The site is arranged with numerous delivery ports, each associated with a specific sorting category, gathering all the parcels of that category. A delivery port can be a bin, a cage cart, or a chute in a dual-layer structure. Robots deliver parcels to delivery ports, serving as the combination of conveyor belts and actuators in conventional sorting systems. Loading stations are located on the periphery of the site, where workers load parcels onto the robots. A top-mounted tray or conveyor actuator enables the robot to load or unload parcels. Figure 1(b) shows a specific model of sorting robot developed by Geekplus, a global technology company specialized in smart logistics and robots. The process for a robot to execute a sorting task is as follows:
@@
    The robot receives a parcel at a loading station, along with the destination information;
@@
    When reaching the target outlet, the robot releases the loaded parcel;
@@
E-commerce order fluctuate during special selling seasons; therefore, warehouse throughput must be designed to accommodate these variations. Unlike conveyor-based sorting systems, the independent and modular nature of robotic operations contributes to the flexibility and scalability of RSS, making it well-suited to dynamic demands of e-commerce (Azadeh et al.,, 2017). During off-hours, warehouse managers can easily configure the control software to change the status of loading stations, the correspondences between outlets and destinations, and the aisle network topology (Xu et al.,, 2022). To leverage the flexibility, robotics companies introduce the innovative Robots-as-a-Service (RaaS) business model, allowing logistics clients to rent robots as needed. RaaS eliminates the high costs associated with purchasing and maintaining robots and peripheral equipment, enabling warehouse managers to adjust the number of rented robots based on demand fluctuations.
@@
Despite these advantages, a research gap remains in investigating the flexibility of RSS and minimizing system costs, as accurately modeling system efficiency is challenging. In robotic systems, robots rely on path planning and coordination algorithms to complete tasks. The problem of finding conflict-free trajectories for all robots is known as the multi-agent path finding (MAPF) problem. The MAPF problem has been proven to be NP-hard (Yu and LaValle,, 2013), which means that finding a relatively good solution can be time-consuming, especially when the scale of the robot system increases, resulting in the instability of RSS efficiency. Shi et al., (2021) analyzed a dataset from the China Post sortation center and found that the congestion effect significantly hampers robot efficiency in RSS, particularly when dealing with a substantial parcel flow. Zou et al., (2021) proposed several closed queuing network (CQN) models to quantitatively analyze the impact of congestion and designed an algorithm to estimate throughput. Validated through numerical experiments, these CQN models accurately reflect the performance of a real case from Deppon Express. However, the models primarily focus on the process of agents transferring between queues rather than on traffic flow and do not fully consider traffic management methods to mitigate conflicts. This aspect makes them less suited for analyzing the efficiency of MAPF solvers, such as the method proposed in this paper. On the other hand, other existing studies often overlooked issues related to robot coordination, relying on the idealized assumption that there are no conflicts or deadlocks. Consequently, their results may become distorted as the actual travel time of robots deviates from the free-flow travel time. Furthermore, these studies do not adequately explore the flexibility of RSS, particularly the ability to adjust the number of robots.
@@
Given the existing gap, this paper aims to address critical challenges in designing an efficient RSS, focusing on three key areas. First, we propose a novel robot traffic management method to enhance system throughput, ensuring stable and efficient operations in multi-robot collaboration settings while overcoming the complexities of robot interactions. Second, we develop a precise model to quantify system efficiency in RSS, enabling accurate throughput estimation to clarify the relationship between system configurations and sorting capacity. Third, we highlight the need for management strategies that leverage RSS flexibility to provide cost-effective responses to fluctuating demands. These contributions aim to improve the performance and scalability of RSS implementations.
@@
To address the aforementioned challenges, we first explore the potential of centralized control strategy for sorting robots. To fully utilize global information while considering robot operational patterns and safety distance regulations, we adapt a network-level control strategy for coordinating autonomous vehicles and propose a collision-free spatio-temporal path planning method for RSS. An efficient heuristic algorithm is proposed to enhance the scalability. Second, building upon this control strategy, we derive an estimation formula for throughput in RSS given system configurations, which serves as an efficiency constraint in the planning problem. Finally, we propose an optimization model aimed at minimizing the initial investment and average operations costs under fluctuating demands. The solution includes layout design and resource allocation recommendations, specifically on how to adjust the number of robots and workers. Experiments provide insights into the cost structure of RSS under varying unit price conditions. To the best of our knowledge, this paper is among the first to incorporate the traffic issues of robots during operations into the configuration design of the RSS system. Our intention is to bridge operational and strategic planning in multi-robot coordination contexts, thereby contributing a comprehensive theoretical framework to support decision-making in such complex and flexible systems.
@@
The remainder of this paper is organized as follows. In section 2, we provide an extensive review of relevant research on RSS systems and conventional sorting systems, highlighting the contributions of this research. Section 3 offers a detailed problem description. In section 4, we introduce an innovative traffic management framework to coordinate multiple robots in a warehouse setting. Based on the proposed framework, section 5 presents a system throughput estimation formula, which provides quantitative measurement of layout configuration’s impact to system efficiency. In section 6, we demonstrate the superiority of our proposed traffic management framework compared with benchmarks, as well as validate the accuracy of throughput estimation formula. To emphasize the impact of traffic issues on RSS, experiments are conducted to show that queueing network model has estimation biases for throughput in certain scenarios. Moving on to section 7, we propose a layout optimization model to minimize the total costs and introduce an efficient solution approach. Section 8 presents a sensitivity analysis to investigate the optimal layout design and the corresponding cost structure under different unit cost, and insights distilled from the results. Finally, section 9 provides a summary of the entire paper.
@@ 2 Literature Review
The cost savings in labor through automation are particularly crucial in the picking process, and increased sorting efficiency enables parcel companies to offer convenient and low-cost same-day delivery services (Dekhne et al.,, 2019). However, the operations research community has scarcely explored the realm of robotized sorting systems, which represent the latest advance of warehouse automation. Among the existing literature, studies mainly focus on the efficiency of small parcel sorting systems in distribution centers. This body of work delves into the comprehensive assessment of transportation efficiency within the robot fleet under various system configurations. In the majority of these studies, robots’ sorting tasks are modeled as services in queuing network models, and the total throughput is the key performance evaluation metrics (Zou and Chen,, 2020; Zi and Gao, 2020a, ; Zi and Gao, 2020b, ; Zou et al.,, 2021; Xu et al.,, 2022). Zi and Gao, 2020a ; Zi and Gao, 2020b assumed that each robot moves independently. Zou and Chen, (2020) introduced a semi-open queuing network (SOQN) model and subsequently improved its theoretical aspects in their follow-up research (Zou et al.,, 2021), simplifying it to the closed queueing model (CQN) by assuming a sufficient high arrival rate of parcels. A queueing model takes into account the queuing behavior of robots at loading stations and outlets, partially addressing the issue of robot occupancy on the aisle. Furthermore, their second study compares how the minimum total costs of the system varied under different network topology and different unit cost of system components, given the target sorting efficiency. Xu et al., (2022) further investigated an RSS with a parcel-to-loading-station assignment mechanism, which could be regarded as a pre-sorting strategy. This study removes the constraint on the number of robots, resulting in an open queuing network (OQN) model. Experimental results demonstrates that the introduction of pre-sorting effectively reduces the average travel distance of the robots but might lead to extra congestion upstream at the same time.
@@
Liu et al., 2019b and Tan et al., (2021) focused on the task assignment problem with the objective of minimizing the sorting makespan. Both adopts a travel time model (closed-form travel time expression, defined by Azadeh et al., (2017)), and the second research formulates a mixed-integer programming model. Compared to queuing network models, travel time models exhibit significant distortion when the number of robots is high, as they allow multiple robots to simultaneously occupy a facility. Boysen et al., (2023) conducted research on order sorting system, which involves additional tasks including Piece-to-Order assignment and Order-to-Collection Point assignment. Therefore, the primary focus of the referred study is not solely on scheduling and control of robots, and the travel time of robots to a specific outlet is assumed to be constant. Shi et al., (2021)’s study, on the other hand, investigates a human-robot hybrid sorting system, in which the scale of robot sorting is fixed, while the manual capacity is adjustable according to the demand. Regrettably, in this context, the considerable flexibility inherent in the robotic system has not been fully harnessed.
@@
Liu et al., 2019b parcel sortation TT, MIP MS, C $\times$
@@
Zi and Gao, 2020a parcel sortation Q T $\times$
Zi and Gao, 2020b parcel sortation Q, MIP T, C $\times$
@@
Tan et al., (2021) parcel sortation TT, MIP MS $\times$
Xu et al., (2022) parcel sortation OQN T $\times$
Boysen et al., (2023) order sortation TT, MIP T $\times$
@@
Throughout the existing research, a limited number of studies consider the floor space cost while others relax the constraints on aisle resources. Furthermore, finding collision-free paths for multiple vehicles has been proved to be an NP-hard problem (Surynek,, 2010), and some state-of-the-art algorithms also struggle to completely eliminate the issue of road congestion resulting from an increasing number of robots. However, except the study by Shi et al., (2021), which employs a fitting method to predict the positive correlation between traffic flow and congestion, most of the studies neglect the mutual influence among multiple robots in aisle occupation for a compact system, leading to an misestimate of traffic capacity. Table 1 presents a comprehensive overview of the literature reviewed on the RSS, highlighting its relevance to our paper.
Some researchers recognized the limitations of queuing models in capturing the impact of aisle layout design on multi-robot systems and chose to address the traffic issues by modeling them as MAPF problems. Wagner and Choset, (2011) proposed a novel M* algorithm that allowed for the planning of a larger number of robot trajectories. However, they faced challenges due to significant memory requirements and time complexity, and could not guarantee a feasible solution. Some researchers designed priority-based algorithms and integrated task assignment to reduce waiting time (Nguyen et al.,, 2019; Liu et al., 2019a, ), sacrificing solution optimality in exchange for stability and computational efficiency. Li et al., (2021) introduced an RHCR algorithm framework to tackle the windowed MAPF problem, providing an efficient solution for managing newly arrived orders in warehouse operations. However, despite their effectiveness, MAPF-based studies have faced challenges in deriving a closed-form expression for throughput estimation, which is crucial for creating reliable and efficient system configurations. Consequently, few studies on multi-robot systems comprehensively address both operational robot management and strategic layout design while considering their interplay.
In summary, existing research on RSS systems lacks sufficient focus on robot traffic management and overall system costs. This study addresses these gaps by first developing a dedicated robot traffic management method for RSS at the operational level. Building on this method, we establish an optimal layout and configuration design model aimed at minimizing overall system costs in the long run, ensuring that the results are practically applicable in real-world warehouses.

### v4

@@ Abstract
In the contemporary logistics industry, automation plays a pivotal role in enhancing production efficiency and expanding industrial scale. In particular, autonomous mobile robots have become integral to modernization efforts in warehouses. One noteworthy application in robotic warehousing is the robotic sorting system (RSS), which is distinguished by its cost-effectiveness, simplicity, scalability, and adaptable throughput control. Previous research on RSS efficiency often assumed an ideal robot management system, ignoring potential traffic delays and assuming constant travel times. To address this gap, we introduce a novel robot traffic management method, named Rhythmic Control for the Sorting Scenario (RC-S), for RSS operations, along with an analytical estimation formula that establishes the quantitative relationship between system performance and configurations. Simulations validate that RC-S reduces average service time by 10.3% compared to the classical cooperative A* algorithm, while also improving throughput and runtime. Based on the performance analysis of RC-S, we develop a layout optimization model that considers system configurations, desired throughput, and costs to minimize expenses and determine the optimal layout. Numerical studies show that facility costs dominate at lower throughput levels, while labor costs prevail at higher throughput levels. Additionally, due to traffic efficiency limitations, an RSS is well-suited for small-scale operations like end-of-supply-chain distribution centers.
@@ 1 Introduction
Improved logistics and delivery services have fueled the rapid growth of e-commerce in the 21st century. The development of stable supply chains, same-day or next-day delivery options, and hassle-free return policies increases consumer confidence in online shopping (ReginaLoBiondo). To handle the increasing volume of online orders, a new generation of warehouses specifically catering to individual customers has become a hot topic for logistics companies. This type of warehouse efficiently meets the demand for small orders with tight delivery schedules through the implementation of automated equipment (BOYSEN2019396). For instance, the robotic mobile fulfillment system (RMFS), which employs a rack-moving mechanism, is widely used in intra-warehouse logistics, such as Amazon’s KIVA systems (wurman2008coordinating).
The main activities performed in a warehouse include: (1) receiving, (2) transfer and put-away, (3) order picking/selection, (4) accumulation/sorting, (5) cross-docking, and (6) shipping (de2007design). In this paper, we focus on the sorting process. Sorting involves categorizing and consolidating parcels according to their order information and shipping destinations. Conventional sorting systems commonly employ conveyor-based sorters , where the actuators move along with the conveyor belt, sequentially passing through each outbound station and releasing the loaded parcels at the appropriate locations (BOYSEN2019796). These systems are highly appreciated for their efficiency and stability, while suffering from the inflexibility and significant space occupation (BOYSEN2023582).
@@
A novel sorting system utilizing mobile robots attracts increasing attention in recent years, namely the robotic sorting system (RSS) (zou2021robotic). Figure 1(a) presents a dual-layer solution of RSS. The site is arranged with numerous delivery ports, each associated with a specific sorting category, gathering all parcels of that category. A delivery port can be a bin, a cage cart, or a chute in a dual-layer structure. Robots deliver parcels to delivery ports, serving as the combination of conveyor belts and actuators in conventional sorting systems. Loading stations are located on the periphery of the site, where workers load parcels onto the robots. A top-mounted tray or conveyor actuator enables the robot to load or unload parcels. Figure 1(b) shows a specific model of sorting robot developed by Geekplus, a global technology company specialized in smart logistics and robots. The process for a robot to execute a sorting task is as follows:
@@
    The robot receives a parcel at a loading station, along with destination information;
@@
    Upon reaching the target outlet, the robot releases the loaded parcel;
@@
E-commerce orders fluctuate during special selling seasons; therefore, warehouse throughput must be designed to accommodate these variations. Unlike conveyor-based sorting systems, the independent and modular nature of robotic operations contributes to the flexibility and scalability of RSS, making it well-suited to the dynamic demands of e-commerce (pub99983). During off-hours, warehouse managers can easily configure the control software to change the status of loading stations, the correspondences between outlets and destinations, and the aisle network topology (XU2022102808). To leverage this flexibility, robotics companies have introduced the innovative Robots-as-a-Service (RaaS) business model, allowing logistics clients to rent robots as needed. RaaS eliminates the high costs associated with purchasing and maintaining robots and peripheral equipment, enabling warehouse managers to adjust the number of rented robots based on demand fluctuations.
@@
Despite these advantages, a research gap remains in investigating the flexibility of RSS and minimizing system costs, as accurately modeling system efficiency is challenging. In robotic systems, robots rely on path planning and coordination algorithms to complete tasks. The problem of finding conflict-free trajectories for all robots is known as the multi-agent path finding (MAPF) problem. The MAPF problem has been proven to be NP-hard (Yu_LaValle_2013), which means that finding a relatively good solution can be time-consuming, especially when the scale of the robot system increases, resulting in instability of RSS efficiency. poms13626 analyzed a dataset from the China Post sortation center and found that the congestion effect significantly hampers robot efficiency in RSS, particularly when dealing with a substantial parcel flow. zou2021robotic proposed several closed queueing network (CQN) models to quantitatively analyze the impact of congestion and designed an algorithm to estimate throughput. Validated through numerical experiments, these CQN models accurately reflect the performance of a real case from Deppon Express. However, the models primarily focus on the process of agents transferring between queues rather than on traffic flow and do not fully consider traffic management methods to mitigate conflicts. This makes them less suitable for analyzing the efficiency of MAPF solvers, such as the method proposed in this paper. On the other hand, other existing studies often overlooked issues related to robot coordination, relying on the idealized assumption that there are no conflicts or deadlocks. Consequently, their results may become distorted as the actual travel time of robots deviates from the free-flow travel time. Furthermore, these studies do not adequately explore the flexibility of RSS, particularly the ability to adjust the number of robots.
@@
Given the existing gap, this paper aims to address critical challenges in designing an efficient RSS, focusing on three key areas. First, we propose a novel robot traffic management method to enhance system throughput, ensuring stable and efficient operations in multi-robot collaboration settings while overcoming the complexities of robot interactions. Second, we develop a precise model to quantify system efficiency in RSS, enabling accurate throughput estimation to reveal the relationship between system configurations and sorting capacity. Third, we highlight the need for management strategies that leverage RSS flexibility to provide cost-effective responses to fluctuating demands. These contributions aim to improve the performance and scalability of RSS implementations.
@@
To address the aforementioned challenges, we first explore the potential of centralized control strategy for sorting robots. To fully utilize global information while considering robot operational patterns and safety distance regulations, we adapt a network-level control strategy for coordinating autonomous vehicles and propose a collision-free spatio-temporal path planning method for RSS. An efficient heuristic algorithm is introduced to enhance the scalability. Second, building upon this control strategy, we derive an estimation formula for throughput in RSS given system configurations, which serves as an efficiency constraint in the planning problem. Finally, we propose an optimization model that aims to minimize initial investment and average operational costs under fluctuating demands. The solution includes layout design and resource allocation recommendations, specifically on how to adjust the number of robots and workers. Experiments provide insights into the cost structure of RSS under varying unit price conditions. To the best of our knowledge, this paper is among the first to incorporate robot traffic issues during operations into RSS configuration design. Our intention is to bridge operational and strategic planning in multi-robot coordination contexts, thereby contributing a comprehensive theoretical framework to support decision-making in such complex and flexible systems.
@@
The remainder of this paper is organized as follows. In Section 2, we provide an extensive review of relevant research on RSS systems and conventional sorting systems, highlighting the contributions of this research. Section 3 offers a detailed problem description. In Section 4, we introduce an innovative traffic management framework to coordinate multiple robots in a warehouse setting. Based on the proposed framework, Section 5 presents a system throughput estimation formula, which quantifies the impact of layout configuration on system efficiency. Section 6 demonstrates the superiority of our proposed traffic management framework compared with benchmark methods and validates the accuracy of the throughput estimation formula. To emphasize the impact of traffic issues on RSS, experiments are conducted to show that queueing network model has estimation biases for throughput in certain scenarios. Moving on to Section 7, we propose a layout optimization model to minimize total costs and introduce an efficient solution approach. Section 8 presents a sensitivity analysis to investigate the optimal layout design and the corresponding cost structure under different unit costs, and insights distilled from the results. Finally, section 9 provides a summary of the entire paper.
@@ 2 Literature Review
The cost savings in labor through automation are particularly crucial in the picking process, and improved sorting efficiency enables parcel companies to offer convenient and low-cost same-day delivery services (dekhne2019automation). However, the operations research community has scarcely explored the realm of robotized sorting systems, which represent the latest advances in warehouse automation. Among the existing literature, studies focus mainly on the efficiency of small parcel sorting systems in distribution centers. This body of work delves into the comprehensive assessment of transportation efficiency within the robot fleet under various system configurations. In the majority of these studies, robot sorting tasks are modeled as services in queuing network models, and total throughput is the key performance evaluation metric (9532245; 9216812; 9102081; zou2021robotic; XU2022102808; fang2025dynamic). 9216812; 9102081 both assumed that each robot moves independently. 9532245 introduced a semi-open queuing network (SOQN) model and subsequently improved its theoretical aspects in their follow-up research (zou2021robotic), simplifying it to the closed queueing model (CQN) by assuming a sufficiently high parcel arrival rate. Their queueing model took into account the queuing behavior of robots at loading stations and outlets, partially addressing the issue of robot occupancy on the aisle. Furthermore, their second study compared how the minimum total costs of the system varied under different network topologies and different unit costs of system components, given the target sorting efficiency. fang2025dynamic further emphasized the importance of dynamic robot routing and destination assignment policies in handling severe heterogeneity and fluctuations in sorting demand. They proposed an SOQN model to analyze network congestion and an MIP model to reassign the destination-outlet mapping in response to drastic changes in sorting demand. XU2022102808 investigated an RSS with a parcel-to-loading-station assignment mechanism, which could be regarded as a pre-sorting strategy. This study removed the constraint on the number of robots, resulting in an open queueing network (OQN) model. Experimental results demonstrated that the introduction of pre-sorting effectively reduces the average travel distance of the robots but might simultaneously lead to additional upstream congestion.
@@
liu2019multi and TAN2021101279 focused on the task assignment problem with the objective of minimizing the sorting makespan. Both adopt a travel time model based on the closed-form travel time expression proposed by pub99983, and the latter formulates a mixed-integer programming model. Compared to queuing network models, travel time models exhibit significant deviation when the number of robots is high, as they allow multiple robots to simultaneously occupy a facility. BOYSEN2023582 conducted research on an order sorting system, which involves additional tasks including Piece-to-Order assignment and Order-to-Collection Point assignment. Therefore, the primary focus of their study is not solely on scheduling and control of robots, and the travel time of robots to a specific outlet is assumed to be constant. poms13626, on the other hand, investigates a human-robot hybrid sorting system, in which the scale of robot sorting is fixed, while the manual capacity is adjustable according to the demand. Regrettably, in this context, the considerable flexibility inherent in the robotic system has not been fully exploited.
@@
liu2019multi parcel sortation TT, MIP MS, C $\times$
@@
9216812 parcel sortation Q T $\times$
9102081 parcel sortation Q, MIP T, C $\times$
@@
TAN2021101279 parcel sortation TT, MIP MS $\times$
XU2022102808 parcel sortation OQN T $\times$
BOYSEN2023582 order sortation TT, MIP T $\times$
[…]
@@
Throughout the existing research, a limited number of studies consider the floor space cost while others relax the constraints on aisle resources. Furthermore, finding collision-free paths for multiple vehicles has been proved to be an NP-hard problem (Surynek_2010), and some state-of-the-art algorithms also struggle to completely eliminate the issue of aisle congestion resulting from an increasing number of robots. However, except for the study by poms13626, which employs a fitting method to predict the positive correlation between traffic flow and congestion, most of the studies neglect the mutual influence among multiple robots in aisle occupation for a compact system, leading to a misestimate of traffic capacity. Table 1 presents a comprehensive overview of the literature reviewed on the RSS, highlighting its relevance to our paper.
Some researchers recognized the limitations of queuing models in capturing the impact of aisle layout design on multi-robot systems and chose to address the traffic issues by modeling them as MAPF problems. wagner6095022 proposed a novel M* algorithm that allowed for the planning of a larger number of robot trajectories. However, they faced challenges due to significant memory requirements and time complexity, and could not guarantee a feasible solution. Some researchers designed priority-based algorithms and integrated them with task assignment to reduce waiting time (nguyen2019generalized; Liu2019path), sacrificing optimality for improved stability and computational efficiency. Li_Tinka2021 introduced an RHCR algorithm framework to tackle the windowed MAPF problem, providing an efficient solution for managing newly arrived orders in warehouse operations. However, despite their effectiveness, MAPF-based studies have faced challenges in deriving a closed-form expression for throughput estimation, which is crucial for creating reliable and efficient system configurations. Consequently, few studies on multi-robot systems comprehensively address both operational robot management and strategic layout design while considering their interplay.
In summary, existing research on RSS systems lacks sufficient focus on robot traffic management and overall system costs. This study addresses these gaps by first developing a dedicated robot traffic management method for RSS at the operational level. Building upon this method, we establish an optimal layout and configuration design model aimed at minimizing overall system costs in the long run, thus enhancing the applicability of the findings to real-world logistics operations.